# Temporal Drift in Privacy Recall: Users Misremember From Verbatim Loss to Gist-Based Overexposure Over Time

### v1

@@ 5. Limitations and Future Work
Temporal drift likely also interacts with some other cues that we have not foregrounded here: thumbnails, location, comments, and algorithmic surfacing signals (i.e., why is this item resurfacing). Field deployments should investigate whether the presence of clear provenance competes with or complements these cues and if copy such as ”Original audience: Friends (May 2024)” changes the social meaning attributed to the re-sharel or its chilling potential. The longer-term social consequences remain an open question: does making prior audiences more salient prompt cultures to reuse with extra care, or does it inadvertently constrain self-presentation (Hogan, 2010; Bazarova and Choi, 2014; Taddicken, 2014; Stutzman et al., 2013; Kokolakis, 2017; Baruh et al., 2017)?
Beyond the reuse within a single application, this is another extremely important frontier. Lack of interoperable privacy-preserving provenance standards implies that context can easily collapse across apps (Sivan-Sevilla and Poudel, 2024). Profiling and reasoning about existing standards such as W3C PROV in terms of the audience claims and use cases would outline safer fallbacks for lossy mappings, thus furnishing designers with workable building blocks (Moreau and Missier, 2013). In some cultures and communities (e.g., stigmatized groups, and adult content creators (Bonifacio et al., 2025)), revealing certain provenance labels is sensitive to whom would be the subject of a research agenda in graduated disclosure for a combined usefulness and safety perspective.
We suggested evaluation involving opt-in instrumentation, log-only simulation, and conservative priors as opposed to personal user studies. This is a feature rather than a limitation once one is dealing with potentially sensitive content and vulnerable populations. Where practical and when ethical review allows, field trials can in fact calibrate priors and test portability with strong privacy guarantees. When experimentation is not viable, we suggest presenting sensitivity bands and focus on interventions that dominate over ranges of $p_{c},p_{o}$ consistent with the literature (Acquisti et al., 2015; Brandimarte et al., 2013; Tversky and Kahneman, 1974; Kahneman, 2011).
Finally, beyond social media, analogous drift appears in document management, messaging, and media vaults (e.g., private galleries (Geeng et al., 2024)). Adapting badges and restoration prompts to those domains, with domain-specific audience ontologies and higher-stakes defaults, is promising. Richer UI exemplars (e.g., mockups vetted with accessibility experts) and open-source pattern libraries could accelerate adoption.

### v2

@@ 5. Limitations and Future Work
Temporal drift likely also interacts with some other cues that we have not foregrounded here: thumbnails, location, comments, and algorithmic surfacing signals (i.e., why is this item resurfacing). Field deployments should investigate whether the presence of clear provenance competes with or complements these cues and if copy such as ”Original audience: Friends (May 2024)” changes the social meaning attributed to the re-sharel or its chilling potential. The longer-term social consequences remain an open question: does making prior audiences more salient prompt cultures to reuse with extra care, or does it inadvertently constrain self-presentation (Hogan2010PresentationOfSelf; Bazarova2014SelfDisclosureJCMC; Taddicken2014JCMCPrivacyParadox; Stutzman2013SilentListeners; Kokolakis2017PrivacyParadox; Baruh2017MediatizedPrivacyParadox)?
Beyond the reuse within a single application, this is another extremely important frontier. Lack of interoperable privacy-preserving provenance standards implies that context can easily collapse across apps (sivanSevilla2024contextualIntegrity). Profiling and reasoning about existing standards such as W3C PROV in terms of the audience claims and use cases would outline safer fallbacks for lossy mappings, thus furnishing designers with workable building blocks (Moreau2013PROVDM). In some cultures and communities (e.g., stigmatized groups, and adult content creators (bonifacio2025onlyfans)), revealing certain provenance labels is sensitive to whom would be the subject of a research agenda in graduated disclosure for a combined usefulness and safety perspective.
We suggested evaluation involving opt-in instrumentation, log-only simulation, and conservative priors as opposed to personal user studies. This is a feature rather than a limitation once one is dealing with potentially sensitive content and vulnerable populations. Where practical and when ethical review allows, field trials can in fact calibrate priors and test portability with strong privacy guarantees. When experimentation is not viable, we suggest presenting sensitivity bands and focus on interventions that dominate over ranges of $p_{c},p_{o}$ consistent with the literature (Acquisti2015PrivacyScience; Brandimarte2013MisplacedConfidences; Tversky1974Judgment; Kahneman2011Thinking).
Finally, beyond social media, analogous drift appears in document management, messaging, and media vaults (e.g., private galleries (Geeng2024Vault)). Adapting badges and restoration prompts to those domains, with domain-specific audience ontologies and higher-stakes defaults, is promising. Richer UI exemplars (e.g., mockups vetted with accessibility experts) and open-source pattern libraries could accelerate adoption.